You are considering a new product launch. The project will cost $820,000, have a 4-year life, and have no salvage value; depreciation is straight-line to zero. Sales are projected at 490 units per year; price per unit will be $18,400, variable cost per unit will be $15,100, and fixed costs will be $830,000 per year. The required return on the project is 14 percent, and the relevant tax rate is 25 percent.

  

a.

The unit sales, variable cost, and fixed cost projections given above are probably accurate to within ±10 percent. What are the upper and lower bounds for these projections? What is the base-case NPV? What are the best-case and worst-case scenarios? (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your NPV answers to 2 decimal places, e.g., 32.16.)


    


    


b.

Calculate the sensitivity of your base-case NPV to changes in fixed costs. (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

c.

What is the accounting break-even level of output for this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

o Accounting Break even part = fixed cost- Sale Pirie - Variable cost 83000 1840-15 Accounting Break wen pant = 251.52 units

6 Sensitivity Let us the base Analysis assume that : Discounted tran Fixe cost wil traghen by 5% not c.f will be Sales = 490Y A wout au ! Salu = 490 -10% = 441 Sale Panie 19400 -- 10% = 11560 Vanable cost = 15100 + 10% = 16610 Fixed cost $3000-tp%.Best Case Nev sales = 490 +10% = 539 Sale Price per unit -$1846 to% = $50,20 Vayable cost lunt = $15100 - 10% = $13,590 FixedBass-case Nor Part. sales 90,16 00 Total variable cost 13,99, 00 Fried Cost Svoor 99 . 20 0 C EBT. Tam enp Not I nome Dey. Neuppy & Land Bounds ay Part base Case - t what Saly 490 ale lun 18400 Varede Cast 15100 Fixed last 830000 withers 10% eppes Bo